# Subhalos are Distributed Anisotropically About Their Hosts

### v1

@@ Abstract
We investigate the distributions of subhalos about their hosts in two suites of zoom-in N-body simulations of halo growth – one suite focused on Milky Way Mass halos ($\sim 10^{12}\mathrm{M}_{\odot}$) and another focused on cluster ($\sim 10^{15}\mathrm{M}_{\odot}$) halos in the Symphony simulation suite. We find, in agreement with previous work on this subject, that subhalos are distributed anisotropically about their host halos. In particular, the positions of subhalos lie preferentially near the major axes of their host halos, possibly implying that satellite galaxies will exhibit a similar alignment. Furthermore, we show that in two-dimensional projection subhalos are more likely to be observed near the halo center (where the central galaxy presumably resides) when the host halo is projected nearly along its major axis. This projection effect is significant. Within a projected radius of $5\%$ of the virial radius of the host halo, the fraction of mass in subhalos is $\sim 44\%$ larger for Milky Way mass halos and as much as $\sim 145\%$ larger for cluster halos when projected along the major axis as compared to the average from a random projection. This result has consequences for many applications including the interpretation of gravitational lenses. Finally, we find that the orbital angular momentum vector of subhalos is aligned with the angular momentum vector of their host halo, indicating that a significant component of a halo’s angular momentum may be carried in its subhalos. This has consequences for galaxy formation models which use host halo angular momentum as a proxy for galaxy momentum.
@@
pubyear: 2022pagerange: Subhalos are Distributed Anisotropically About Their HostsLABEL:lastpage
@@ 1 Introduction
The contemporary galaxy formation paradigm holds that galaxies form within the potential wells supplied by halos of dark matter (White & Rees, 1978; Blumenthal et al., 1984). In the standard $\Lambda$CDM (cold dark matter with a cosmological constant) model of the universe, dark matter halos form when perturbations in the initial matter field collapse into gravitationally-bound objects. These structures grow by the accretion of mass from surrounding dark matter and through mergers with other halos; smaller halos that are not entirely disrupted upon merging with a larger halo are known as subhalos (see e.g., Kauffmann et al. 1993; Klypin et al. 1999; Moore et al. 1999; Taylor & Babul 2002; Zentner & Bullock 2003; Taylor & Babul 2004; Zentner et al. 2005a; Bullock 2010; Jiang & van den Bosch 2016; van den Bosch & Jiang 2016).
$\Lambda$CDM predicts the existence of subhalos on all mass scales. Some alternative models for dark matter predict differing amounts of substructure. Warm dark matter (WDM) has become the benchmark example of an alternative to CDM that predicts less substructure in general and fewer subhalos in particular (e.g., Colín et al., 2000; Bode et al., 2001; Lovell et al., 2014). However, various alternatives to CDM make distinct predictions for the structure and abundance of dark matter, including self-interacting dark matter (e.g., Tulin & Yu, 2018), fuzzy dark matter (e.g., Hui, 2021), and decaying dark matter (e.g., Peter & Benson, 2010; Wang et al., 2013, 2014). Consequently, the observation (or non-observation) of subhalos within larger host halos can be used as a test of dark matter.
@@
A significant body of work predicts that subhalos in CDM simulations of structure formation should be distibuted anisotropically about their host halos. Host dark matter halos are ellipsoidal in shape (Barnes & Efstathiou, 1987; Dubinski & Carlberg, 1991; Warren et al., 1992; Dubinski, 1994; Tormen, 1997; Thomas et al., 1998; Jing & Suto, 2002; Springel et al., 2004; Hopkins et al., 2005), and in particular, halos tend to be prolate ellipsoids. Zentner et al. 2005b found preferential alignments between the overall mass distribution of Milky Way-mass host halos and the spatial distribution of their subhalos, and this was confirmed in galaxy formation simulations shortly thereafter (Libeskind et al., 2007). More recently, Karp et al. (2023) find an overdensity of satellite galaxies along the major axis of their central galaxy in the IllustrisTNG simulation. It is believed that these phenomena are the result of halos preferentially merging along filaments in the large-scale structure (Wang et al., 2005; Zentner et al., 2005b; Libeskind et al., 2007; Libeskind et al., 2011; Libeskind et al., 2015; Shi et al., 2015; Kang & Wang, 2015; Shao et al., 2018; Morinaga & Ishiyama, 2020). These filaments trace out preferred directions about halos in the large scale structure. Several simulations have found that alignments between neighboring dark matter halos tend to take place along the direction of the longest axes (the major axes) of the halos. Faltenbacher et al. 2002 and Hopkins et al. 2005 detect alignments between the major axes of cluster-sized halos out to separations of approximately 100 $\textrm{h}^{-1}$ Mpc and 30 $\textrm{h}^{-1}$ Mpc, respectively. Using the IllustrisTNG simulation, Rodriguez et al. (2024) reports alignments out to 10 $\textrm{h}^{-1}$ Mpc for both the stellar and dark matter components of galaxies.
@@
The prediction that subhalos are disributed anisotropically about their hosts and that they are aligned preferentially with the non-spherical dark matter distribution of their hosts has some observational support from a variety of analyses of galaxies from surveys such as the Sloan Digital Sky Survey (SDSS). Using a sample of galaxy groups from SDSS DR4, Brainerd (2005) and Yang et al. (2006) showed that satellites tend to be aligned with the major axis of the central galaxy in the group. Azzaro et al. (2007) reached qualitatively similar conclusions for the SDSS DR4 sample using complementary methods (see also Azzaro et al., 2006). Wang et al. (2018) and Rodriguez et al. (2022) performed a similar analyses using SDSS DR7 and SDSS DR16 data respectively and found that this alignment exists primarily between red central galaxies and their red satellites. More recently, Samuels & Brainerd (2023) showed that satellite distributions tend to be lopsided with satellites preferentially residing on one side of their host. Similar results have been observed in cluster lensing analyses. Using SDSS Evans & Bridle (2009), determined that dark matter halos are preferentially aligned in the same direction as indicated by the spatial distribution of cluster galaxies using cluster lensing. Gonzalez et al. (2021) drew similar conclusions from a galaxy lensing analysis of a number of surveys.
@@ 2.1 Milky Way and Cluster Mass Zoom-In Simulations
We use two sets of zoom-in cosmological simulations from the Symphony suite, each representing a different halo mass range. The first set of simulations, which were first presented in (Mao et al., 2015), contains high-resolution zoom-in simulations of 45 Milky Way-mass halos from a c125-2048 parent box run with L-GADGET (see Becker 2015). Halo masses in this catalog are $M_{\rm vir}=10^{12.09\pm 0.02}\ \mathrm{M}_{\odot}$. The second set contains zoom-in simulations of 96 cluster-mass halos within the mass range $10^{14.96\pm 0.03}M_{\odot}$ first presented in (Wu et al., 2013). We exclude one halo from the cluster-mass set due to a potential mislabeling between host halo and most massive subhalo.
@@ 2.2 Halo Identification
  • subhalo only: These are particles associated with at least one subhalo within the host halo, as identified by $\tt{ROCKSTAR}$.
@@ 3.2 Halo and Subhalo Angular Momentum
where $m_{n}$ is the mass of the $n^{\rm th}$ particle. The total angular momentum, $\vec{J}_{\mathrm{Tot}}$, includes the angular momentum associated with both host halo and subhalo particles and is computed by summing over the total halo particle set. The host angular momentum, $\vec{J}_{\mathrm{Host}}$, is calculated using Equation 2 by summing over the host only particle set, while the angular momentum carried by the subhalos, $\vec{J}_{\mathrm{Sub}}$, is defined to be the difference $\vec{J}_{\mathrm{Sub}}$ = $\vec{J}_{\mathrm{Tot}}$ $\vec{J}_{\mathrm{Host}}$.
@@ 3.3 The Subhalo Mass Distribution
    Define a cylinder, with a central axis that points along the halo major axis and intersects the halo center, of radius $r<r_{\rm{VIR}}$, and length such that it extends to the bounds of the halo.
@@
  3. 3.
    Rotate the cylinder about the halo center by some random zenith angle, $\theta$, and random azimuthal angle, $\phi$, so that it is projected along a random axis and repeat (ii).
@@
    Repeat (iii) for 100 more random axes.
Figure 1: Diagram showing how cylinders are defined to calculate projected mass as described in Section 3.3. The light pink ellipsoid represents the host halo and the small dark pink circles represent subhalos. The left image shows a cylinder of radius $r$ oriented near the major axis. The cylinder is then rotated to some random orientation as shown in the right image. When comparing the left and right images, we can see how the number of subhalos within cylinders of the same size can vary depending on orientation.
This procedure is repeated for each halo of the Milky Way-mass and cluster-mass halos separately for cylinders with radii equal to 5% and 10% of their host virial radii. These radii were selected because they correspond roughly to the size of the Einstein Radius, where strong lensing takes place. By definition, a radius of $r=0$ encompasses none of the halo mass, whereas a cylinder of radius $r=r_{\rm{VIR}}$ encompasses the entire halo. Once completed, for both cluster-mass and Milky Way-mass halos, we will have a distribution of subhalo mass fractions within a cylinder of radius $r$ as a function of the angle $\theta$ between the cylinder’s axis and the halo’s major axis. For each cylinder radius, we normalize individual subhalo mass fractions by the median mass fraction across all angles and halos (Milky Way-mass and cluster-mass separately). We then bin the data based on values of $\theta$ and take the median value per bin. We use 23 bins for the Milky Way-mass data and 50 for the cluster-mass data. Errors are estimated by building 1000 mock samples of median subhalo mass fractions. Each mock sample contains the same number of bins as its corresponding data (23 or 50). The errors are then the $\rm{16^{th}}$ and $\rm{84^{th}}$ percentiles of these mock samples.
@@ 4 Results I: Host Halo Alignments
As discussed above, we expect that the orientation of $\vec{J}_{\mathrm{Host}}$ with respect to host halo major axes is not isotropic and will be preferentially aligned with the minor axes, and perpendicular to the major axes. Our data is suggestive of an orthogonal alignment between the major axis and angular momentum vector, $\vec{J}_{\mathrm{Host}}$, as expected. In both panels of Fig. 2, we see an excess of halos for which the angle between the angular momentum vector and the major axis, $|\cos(\theta)|$, is small, which corresponds to a perpendicular alignment. The Kolmogorov-Smirnov (KS) probability that these angular distributions are selected from an isotropic distribution is $\sim$ 9 $\times$ 10-7 for Milky Way-mass halos and $\sim$ 3 $\times$ 10-5 for cluster- mass halos, indicating they are far from an isotropic distribution. This alignment has previously been detected in works such as Shaw et al. (2006); Knebe et al. (2008); Deason et al. (2011); Hoffmann et al. (2014); Kiessling et al. (2015).
We can also quantify the alignment between host halo angular momentum vector and host major axis with the median angular separation between the two. For Milky Way-mass halos, the median angular separation is $\theta$ = $76\pm 2.4^{\circ}$ or $|\cos(\theta)|$ = 0.24 $\pm$ 0.04. For Cluster-mass halos, the median angular separation is $\theta$ = $75\pm 2.2^{\circ}$ or cos($\theta$) = 0.25 $\pm$ 0.03. Errors are calculated by bootstrapping. For each set of angles, we sample the angles with replacement to create a mock sample of the same size and calculate its median. To get the uncertainty we repeat this 1000 times and calculate the standard deviation of the set of 1000 medians. For an isotropic distribution of angles, we would expect a median of 60, however, we are measuring angles that are approximately 15 greater. Angles significantly greater than 60 are indicative of perpendicular alignment which is consistent with the discussion in the previous paragraph that a halo’s angular momentum vector is preferentially perpendicular to its major axis.
Figure 2: The cumulative fraction of halos with angle less than $\rm cos(\theta)$ between the host halo angular momentum vector, $\vec{J}_{\mathrm{Host}}$ and its major axis $\vec{A}_{\mathrm{Host}}$ as a function of $\rm cos(\theta)$. The solid line is the result from the simulation data, the dashed line and the filled area are the CDF for an isotropic distribution and the area within the 14th and 86th percentiles. The percentiles are computed from random samples that are the same size as the simulation samples. Results for Milky Way-mass halos are shown in (a) and cluster-mass in (b). The Kolmogorov-Smirnov (KS) probability that the angular distributions are selected from an isotropic distribution, $P_{\rm KS}$, are $\sim$ 9 $\times$ 10-7 for Milky Way-mass and $\sim$ 3 $\times$ 10-5 for cluster-mass.
@@ 5.1 Principle Axis and Angular Momentum Alignments of Subhalos and Their Hosts
We first discuss alignments between the major axis of the distribution of all subhalos belonging to a particular host halo and the major axis of said host halo. The principal axes of the subhalo component are defined as the eigenvectors of the inertia tensor corresponding to subhalo particles only. We calculated the angular separation between the major axis of the host, $\vec{A}_{\mathrm{Host}}$, and that of the subhalos, $\vec{A}_{\mathrm{Sub}}$. Previous works have shown that this result can vary as a function of the radius at which the calculation is performed (Schneider et al., 2012b). In this work, we chose to use major axis as defined up to the full virial radius of the halo. In Fig. 3, we show the cumulative fraction of halos with angle $<|\rm cos(\theta)|$ between $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ as a function of $|\rm cos(\theta)|$. The left plot (3(a)) shows results for Milky Way-mass halos and the right (3(b)) for cluster-mass halos.
@@
In both panels of Fig. 3 we see that there is a deficit of halos at smaller values of $|\rm cos(\theta)|$, indicating that the angular separation between the major axis of the host halo matter distribution and the major axis of the subhalo distribution is smaller than it would be given an isotropic distribution. This implies, perhaps unsurprisingly, that $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ are preferentially aligned. In other words, the distribution of subhalos is aligned with the distribution of host halo mass. The KS probability that the result for Milky Way-mass halos (3(a)) is selected from an isotropic distribution is $P_{\rm KS}$ is $\sim 6\times 10^{-\rm 15}$. For cluster-mass halos (3(b)) this probability is $\sim 2\times 10^{-23}$.
@@
We can also quantify the alignment between $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ with the median of their angular separation. For Milky Way-mass halos, the median angle between the host halo major axis the subhalo component major axis is $\theta$ = $27\pm 7.7^{\circ}$ or $|\cos(\theta)|$ = 0.89 $\pm$ 0.04. For Cluster-mass halos, this median angle is $\theta$ = $15\pm 2.6^{\circ}$ or $|\cos(\theta)|$ = 0.97 $\pm$ 0.01. Errors are calculated with the same bootstrap method as Section 4. For an isotropic distribution of angles (spanning from 0 to 90), we would expect a median of 60, however, we measure angular separations that are approximately 35- 45 smaller. A smaller angle corresponds to greater alignment between major axes and is consistent with the discussion in the previous paragraph that subhalos will reside near their host major axis.
@@
The alignment we find between $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ is not unexpected, a factor to which we have already alluded. Within the model of hierarchical structure growth, subhalos merge with host halos in a correlated manner along filaments (e.g., Plionis & Basilakos, 2002; Jing & Suto, 2002; Bailin & Steinmetz, 2005; Faltenbacher et al., 2005; Zentner et al., 2005b; Shaw et al., 2006; Libeskind et al., 2007; Welker et al., 2014). In a recent work, Han et al. (2023) showed that $\sim 68\%$ of accreted subhalos enter their host halo through $\sim 38\%$ of the halo surface area defined at the virial radius. This results in alignment between the large-scale filamentary structure and that of a host halo (e.g., Jing & Suto, 2002; Bailin & Steinmetz, 2005; Faltenbacher et al., 2005; Zentner et al., 2005b). In turn, this causes the shape and orientation of halos to be ellipsoids with their longest axes pointed along the direction of the filament. Consequently, subhalos are preferentially located near this longest axis, the major axis, and this is the alignment that we are detecting.
@@
Comparing 3(a) with 3(b), it is evident that the alignment between $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ is more pronounced in cluster-mass halos than in Milky Way-mass halos. This difference may be attributed to the later formation times and larger subhalo abundances of cluster-mass halos, which generally have experienced significantly more recent mergers (Lacey & Cole, 1994; Wechsler et al., 2002; Zentner et al., 2005a). Consequently, this means that the major axis of cluster-mass halos has been more recently influenced by the influx of mergers along neighboring filaments. Collectively, these effects lead to subhalos being positioned more closely to the major axis of cluster-mass halos. Such a mass dependence on the strength of alignments has been proposed before on neighboring halos in works such as Jing (2002); Li et al. (2013); Schneider et al. (2012a); Xia et al. (2017). A follow-up analysis would be necessary in order to confirm this interpretation.

### v2

# Subhalos are Anisotropically Distributed and Aligned with the Smooth Matter Distribution of Their Host Halos

@@ Abstract
We investigate the distributions of subhalos about their hosts in two suites of zoom-in N-body simulations of halo growth – one suite focused on Milky Way-Mass halos ($\sim 10^{12}\mathrm{M}_{\odot}$) and another focused on cluster-mass halos ($\sim 10^{15}\mathrm{M}_{\odot}$) in the Symphony simulation suite. We find, in agreement with previous work on this subject, that subhalos are distributed anisotropically about their host halos. In particular, the positions of subhalos lie preferentially near the major axes of their host halos, possibly implying that satellite galaxies will exhibit a similar alignment. Furthermore, we show that in two-dimensional projection subhalos are more likely to be observed near the halo center (where the central galaxy presumably resides) when the host halo is projected nearly along its major axis. This projection effect is significant. Within projected radii of a few percent of the virial radius of the host halo, the fraction of mass in subhalos is $\sim 175\%$ larger for Milky Way mass halos and as much as $\sim 195\%$ larger for cluster halos when projected along the major axis as compared to the average from a random projection. This result has consequences for many applications including the interpretation of gravitational lenses. Finally, we find that the orbital angular momentum vector of subhalos is aligned with the angular momentum vector of their host halo, indicating that a significant component of a halo’s angular momentum may be carried in its subhalos. This has consequences for galaxy formation models which use host halo angular momentum as a proxy for galaxy momentum.
@@
pubyear: 2022pagerange: Subhalos are Anisotropically Distributed and Aligned with the Smooth Matter Distribution of Their Host HalosReferences
@@ 1 Introduction
[…]
@@
A significant body of work predicts that subhalos in CDM simulations of structure formation should be distibuted anisotropically about their host halos. Host dark matter halos are ellipsoidal in shape (Barnes & Efstathiou, 1987; Dubinski & Carlberg, 1991; Warren et al., 1992; Dubinski, 1994; Tormen, 1997; Thomas et al., 1998; Jing & Suto, 2002; Springel et al., 2004; Hopkins et al., 2005), and in particular, halos tend to be prolate ellipsoids. Zentner et al. 2005b found preferential alignments between the overall mass distribution of Milky Way-mass host halos and the spatial distribution of their subhalos, and this was confirmed in galaxy formation simulations shortly thereafter (Libeskind et al., 2007). More recently, Karp et al. (2023) find an overdensity of satellite galaxies along the major axis of their central galaxy in the IllustrisTNG simulation. It is believed that these phenomena are the result of halos preferentially merging along filaments in the large-scale structure (Wang et al., 2005; Zentner et al., 2005b; Libeskind et al., 2007, 2011, 2015; Shi et al., 2015; Kang & Wang, 2015; Shao et al., 2018; Morinaga & Ishiyama, 2020). These filaments trace out preferred directions about halos in the large scale structure. Several simulations have found that alignments between neighboring dark matter halos tend to take place along the direction of the longest axes (the major axes) of the halos. Faltenbacher et al. 2002 and Hopkins et al. 2005 detect alignments between the major axes of cluster-sized halos out to separations of approximately 100 $\textrm{h}^{-1}$ Mpc and 30 $\textrm{h}^{-1}$ Mpc, respectively. Using the IllustrisTNG simulation, Rodriguez et al. (2024) reports alignments out to 10 $\textrm{h}^{-1}$ Mpc for both the stellar and dark matter components of galaxies.
@@
The prediction that subhalos are disributed anisotropically about their hosts and that they are aligned preferentially with the non-spherical dark matter distribution of their hosts has some observational support from a variety of analyses of galaxies from surveys such as the Sloan Digital Sky Survey (SDSS). Using a sample of galaxy groups from SDSS DR4, Brainerd (2005) and Yang et al. (2006) showed that satellites tend to be aligned with the major axis of the central galaxy in the group. Azzaro et al. (2007) reached qualitatively similar conclusions for the SDSS DR4 sample using complementary methods (see also Azzaro et al., 2006). Wang et al. (2018) and Rodriguez et al. (2022) performed a similar analyses using SDSS DR7 and SDSS DR16 data respectively and found that this alignment exists primarily between red central galaxies and their red satellites. More recently, Samuels & Brainerd (2023) showed that satellite distributions tend to be lopsided with satellites preferentially residing on one side of their host in agreement with Libeskind et al. (2016) and Brainerd & Samuels (2021). Similar results have been observed in cluster lensing analyses. Using SDSS Evans & Bridle (2009), determined that dark matter halos are preferentially aligned in the same direction as indicated by the spatial distribution of cluster galaxies using cluster lensing. Gonzalez et al. (2021) drew similar conclusions from a galaxy lensing analysis of a number of surveys.
@@ 2.1 Milky Way and Cluster Mass Zoom-In Simulations
[…]
@@ 2.2 Halo Identification
  • subhalo only: These are particles associated with at least one subhalo within the host halo, as identified by $\tt{ROCKSTAR}$. Subhalos within the symphony simulation are defined such that they contain a minimum of 300 particles at z=0.
@@ 3.1 The Major Axis of the Host Halo
Figure 1: Distributions of halo shape as measured by the ratio of the halo minor axis to major axis ($c/a$) and the semi-major to major axis ($b/a$). The filled histograms correspond to c/a and hollow histograms to $b/a$. Milky Way-mass halos are shown in panel ($a$) on the left and cluster-mass halos on right in panel ($b$). Cluster-mass halos are generally less spherical.
@@ 3.2 Halo and Subhalo Angular Momentum
where $m_{n}$ is the mass of the $n^{\rm th}$ particle. The total angular momentum, $\vec{J}_{\mathrm{Tot}}$, includes the angular momentum associated with both host halo and subhalo particles and is computed by summing over the total halo particle set. The host angular momentum, $\vec{J}_{\mathrm{Host}}$, is calculated using Equation 2 by summing over the host only particle set, while the angular momentum carried by the subhalos, $\vec{J}_{\mathrm{Sub}}$, is defined to be the difference $\vec{J}_{\mathrm{Sub}}$ = $\vec{J}_{\mathrm{Tot}}$ - $\vec{J}_{\mathrm{Host}}$.
@@ 3.3 The Subhalo Mass Distribution
    Define a set of axes along which to project host halo and subhalo mass. The axes are centered on the host halo and spaced isotropically over the surface of a sphere.
@@
    Define a cylinder, with a central axis that points along the halo major axis and intersects the halo center such that the circular face of the cylinder has a radius $r<r_{\rm{VIR}}$ and the length of the cylinder extends to the virial radius of the halo.
  3. 3.
@@
    Rotate the cylinder about the halo center so that it points along the subsequent axis in the predefined set of projection axes.
  5. 5.
    Calculate the angular separation, $\theta$, between the major axis and the new axis of projection.
  6. 6.
    Repeat the mass measurement from step (iii).
  7. 7.
    Repeat the process of rotating and measuring the subhalo mass [steps (iii)-(vi)] within each rotated cylinder for all the axes in the predefined set of axis directions from step (i).
Figure 2: Diagram showing how cylinders are defined to calculate projected mass as described in Section $3.$ The dotted pink circle represents the region enclosed within the virial radius of the halo. The light pink ellipsoid represents the shape of the host halo even though all particles within the virial radius are used, and the filled in dark pink circles represent subhalos. The left image shows a cylinder of radius $r$ oriented near the major axis. The cylinder is then rotated to some random orientation as shown in the right image. When comparing the left and right images, we can see how the number of subhalos within cylinders of the same size can vary depending on orientation.
This procedure is repeated for each halo of the Milky Way-mass and cluster-mass halos separately for cylinders with logorithmically-spaced radii from 0.001 to 1 times their host virial radius. By definition, a radius of $r=0$ encompasses none of the halo mass, whereas a cylinder of radius $r=r_{\rm{VIR}}$ encompasses the entire halo. Once completed, for both cluster-mass and Milky Way-mass halos, we will have a distribution of subhalo mass fractions within a cylinder of radius $r$ as a function of the angle $\theta$ between the cylinder’s axis and the halo’s major axis.
@@ 4 Results I: Host Halo Alignments
As discussed above, we expect that the orientation of $\vec{J}_{\mathrm{Host}}$ with respect to host halo major axes is not isotropic and will be preferentially aligned with the minor axes, and perpendicular to the major axes. Our data is suggestive of an orthogonal alignment between the major axis and angular momentum vector, $\vec{J}_{\mathrm{Host}}$, as expected. In both panels of Fig. 3, we see an excess of halos for which the angle between the angular momentum vector and the major axis, $|\cos(\theta)|$, is small, which corresponds to a perpendicular alignment. The Kolmogorov-Smirnov (KS) probability that these angular distributions are selected from an isotropic distribution is $\sim$ 4 $\times$ 10-5 for Milky Way-mass halos and $\sim$ 3 $\times$ 10-5 for cluster-mass halos, indicating that an isotropic, uncorrelated distribution of halo angular momenta and principle axes is excluded by the data. This alignment has previously been detected in works such as Shaw et al. (2006); Knebe et al. (2008); Deason et al. (2011); Hoffmann et al. (2014); Kiessling et al. (2015).
We can also quantify the alignment between host halo angular momentum vector and host major axis with the median angular separation between the two. For Milky Way-mass halos, the median angular separation is $\theta$ = $77\pm 3.0^{\circ}(17^{\circ})$ or $|\cos(\theta)|$ = 0.22 $\pm$ 0.05 (0.24). The first set of error values are the errors on the medians while the values in the parentheses correspond to the dispersions in the distribution of angles. For Cluster-mass halos, the median angular separation is $\theta$ = $75\pm 2^{\circ}(16^{\circ})$ or cos($\theta$) = 0.25 $\pm$ 0.04 (0.24). Errors are calculated by bootstrapping. For each set of angles, we sample the angles with replacement to create a mock sample of the same size and calculate its median. To get the uncertainty we repeat this 1000 times and calculate the standard deviation of the set of 1000 medians. For an isotropic distribution of angles, we would expect a median of 60, however, we are measuring angles that are approximately 15 greater. Angles significantly greater than 60 are indicative of perpendicular alignment which is consistent with the discussion in the previous paragraph that a halo’s angular momentum vector is preferentially perpendicular to its major axis.
Figure 3: The cumulative fraction of halos with angle less than $\rm cos(\theta)$ between the host halo angular momentum vector, $\vec{J}_{\mathrm{Host}}$ and its major axis $\vec{A}_{\mathrm{Host}}$ as a function of $\rm cos(\theta)$. The solid line is the result from the simulation data, the dashed line and the filled area are the CDF for an isotropic distribution and the area within the 14th and 86th percentiles. The percentiles are computed from random samples that are the same size as the simulation samples. Results for Milky Way-mass halos are shown in (a) and cluster-mass in (b).
@@ 5.1 Principle Axis and Angular Momentum Alignments of Subhalos and Their Hosts
We first discuss alignments between the major axis of the distribution of all subhalos belonging to a particular host halo and the major axis of said host halo. The principal axes of the subhalo component are defined as the eigenvectors of the inertia tensor corresponding to subhalo particles only. We calculated the angular separation between the major axis of the host, $\vec{A}_{\mathrm{Host}}$, and that of the subhalos, $\vec{A}_{\mathrm{Sub}}$. Previous works have shown that this result can vary as a function of the radius at which the calculation is performed (Schneider et al., 2012b). In this work, we chose to use major axis as defined up to the full virial radius of the halo. In Fig. 4, we show the cumulative fraction of halos with angle $|\rm cos(\theta)|$ between $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ as a function of $|\rm cos(\theta)|$. The left plot (4(a)) shows results for Milky Way-mass halos and the right (4(b)) for cluster-mass halos.
@@
In both panels of Fig. 4 we see that there is a deficit of halos at smaller values of $|\rm cos(\theta)|$, indicating that the angular separation between the major axis of the host halo matter distribution and the major axis of the subhalo distribution is smaller than it would be given an isotropic distribution. This implies, perhaps unsurprisingly, that $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ are preferentially aligned. In other words, the distribution of subhalos is aligned with the distribution of host halo mass. The KS probability that the result for Milky Way-mass halos (4(a)) is selected from an isotropic distribution is $P_{\rm KS}$ is $\sim 3\times 10^{-\rm 9}$. For cluster-mass halos (4(b)) this probability is $\sim 7\times 10^{-10}$.
@@
For Milky Way-mass halos, the median angle between the host halo major axis the subhalo component major axis is $\theta$ = $30\pm 4.0^{\circ}(22^{\circ})$ or $|\cos(\theta)|$ = 0.86 $\pm$ 0.04 (0.26). For Cluster-mass halos, this median angle is $\theta$ = $16\pm 3^{\circ}(23^{\circ})$ or $|\cos(\theta)|$ = 0.96 $\pm$ 0.01 (0.24). Errors are calculated with the same bootstrap method as Section 4. For an isotropic distribution of angles (spanning from 0 to 90), we would expect a median of 60; however, we measure angular separations that are approximately 30- 45 smaller. This is consistent with the discussion in the previous paragraph that subhalos preferentially reside closer to the major axis of their host halo.
@@
The alignment we find between $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ is not unexpected, a fact to which we have already alluded. Within the model of hierarchical structure growth, subhalos merge with host halos in a correlated manner along filaments (e.g., Plionis & Basilakos, 2002; Jing & Suto, 2002; Bailin & Steinmetz, 2005; Faltenbacher et al., 2005; Zentner et al., 2005b; Shaw et al., 2006; Libeskind et al., 2007; Welker et al., 2014). In a recent work, Han et al. (2023) showed that $\sim 68\%$ of accreted subhalos enter their host halo through $\sim 38\%$ of the halo surface area defined at the virial radius. This results in alignment between the large-scale filamentary structure and that of a host halo (e.g., Jing & Suto, 2002; Bailin & Steinmetz, 2005; Faltenbacher et al., 2005; Zentner et al., 2005b). In turn, this causes the shape and orientation of halos to be ellipsoids with their longest axes pointed along the direction of the filament. Consequently, subhalos are preferentially located near this longest axis, the major axis, and this is the alignment that we detect.
@@
Comparing 4(a) with 4(b), it is evident that the alignment between $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ is more pronounced in cluster-mass halos than in Milky Way-mass halos. This difference may be attributed to the later formation times and larger subhalo abundances of cluster-mass halos, which generally have experienced significantly more recent mergers (Lacey & Cole, 1994; Wechsler et al., 2002; Zentner et al., 2005a). Consequently, this means that the major axes of cluster-mass halos have been more recently influenced by the influx of mergers along neighboring filaments. Collectively, these effects lead to subhalos being positioned more closely to the major axis of cluster-mass halos. Such a mass dependence on the strength of alignments has been proposed before for neighboring halos in works such as Jing (2002); Li et al. (2013); Schneider et al. (2012a); Xia et al. (2017). A follow-up analysis would be necessary in order to confirm this interpretation.